# AeroGPT: Leveraging Large-Scale Audio Model for Aero-Engine Bearing Fault Diagnosis

### v1

@@ II-B Audio Knowledge Acquisition
The final refinement stage implements a Direct Preference Optimization (DPO) approach. This technique uses a comparative scoring mechanism where alternative responses to the same audio input are evaluated, with scores indicating their relative quality (e.g., Answer 1 with score 8.5 vs. Answer 2 with score 2.0). The optimization objective is formulated as:
@@
where $\mathcal{P}_{\theta}$ represents the system being optimized, $\mathcal{P}_{ref}$ is a reference system, $y_{w}$ and $y_{l}$ are the preferred and less preferred responses respectively, $\sigma$ is the sigmoid function, and $\beta$ is a hyperparameter controlling optimization strength. This comparative approach ensures the system’s acoustic interpretations align with expert assessments, particularly for challenging or ambiguous acoustic patterns.
The culmination of these three stages produces a large-scale audio foundation model with sophisticated acoustic pattern recognition capabilities. This audio knowledge provides the foundation for AeroGPT’s ability to interpret vibration signals from aero-engine bearings as acoustic phenomena, enabling more accurate and nuanced fault diagnosis than conventional approaches.
II-C LoRA-based Domain Knowledge Adaptation
@@
The adaptation process operates on the principle that domain-specific knowledge can be effectively integrated through low-rank decomposition of weight updates. Rather than modifying the entire weight matrices of the audio processing system, which would require enormous computational resources, AeroGPT freezes the pre-trained weight matrices and introduces trainable low-rank adaptations. For a pre-trained weight matrix $\mathbf{W}_{0}\in\mathbb{R}^{d\times k}$ in the model, the adaptation is accomplished through:
$\mathbf{W}=\mathbf{W}_{0}+\Delta\mathbf{W}=\mathbf{W}_{0}+\mathbf{B}\mathbf{A}$ (3)
where $\mathbf{B}\in\mathbb{R}^{d\times r}$ and $\mathbf{A}\in\mathbb{R}^{r\times k}$ are low-rank matrices with rank $r\ll\min(d,k)$. This factorization significantly reduces the number of trainable parameters from $d\times k$ to $r\times(d+k)$.
During the adaptation process, the original pre-trained weight matrix $\mathbf{W}_{0}$ remains frozen, while only the low-rank matrices $\mathbf{A}$ and $\mathbf{B}$ are updated. The forward pass computation is modified as:
$\mathbf{h}=\mathbf{W}_{0}\mathbf{x}+\Delta\mathbf{W}\mathbf{x}=\mathbf{W}_{0}\mathbf{x}+\mathbf{B}\mathbf{A}\mathbf{x}$ (4)
where $\mathbf{x}$ represents the input to the layer and $\mathbf{h}$ is the output. To facilitate stable training, matrix $\mathbf{A}$ is typically initialized with random Gaussian values, while matrix $\mathbf{B}$ is initialized with zeros, ensuring that at the beginning of training, $\Delta\mathbf{W}=\mathbf{B}\mathbf{A}=0$.
In practice, a scaling factor $\alpha$ is also introduced to control the magnitude of the adaptation:
@@
Here, $\alpha$ is a hyperparameter that can be set equal to the rank $r$ or tuned separately. This scaling factor helps stabilize training, especially when experimenting with different rank values.
The LoRA adaptation technique employed in AeroGPT offers unparalleled parameter efficiency with over 99% parameter reduction compared to full fine-tuning approaches. This efficiency translates to significant reductions in VRAM usage, from over 100GB for full fine-tuning to approximately 17GB with LoRA, making the fine-tuning process feasible on widely available hardware. This is particularly advantageous in aerospace applications, where computational resources may be limited.
In our framework, LoRA-based adaption is strategically applied at multiple points. Within the audio encoder, LoRA layers modify the processing of vibration signals, enabling the system to recognize the unique acoustic signatures of various bearing faults. For the language model component, adaptation points are established at the self-attention mechanisms (specifically in query, key, and value projections) and feed-forward networks. Through this knowledge adaptation process, AeroGPT is able to bridge the gap between general audio understanding and the specific characteristics of aero-engine bearing vibrations, establishing a robust foundation for effective fault diagnosis.
II-D Vibration Signal Alignment (VSA)
The Vibration Signal Alignment (VSA) mechanism is designed to bridge the fundamental gap between general audio knowledge and the domain-specific characteristics of aero-engine bearing vibrations. This is achieved through a alignment process that preserves the rich hierarchical representations inherent in large-scale audio models while adapting them to the specialized patterns of mechanical vibrations.
@@ II-D Vibration Signal Alignment (VSA)
This training paradigm enables the model to learn contextual relationships between vibration patterns and diagnostic interpretations in an end-to-end manner while preserving the linguistic capabilities of the underlying large language model, allowing for the generation interpretable diagnostic outputs rather than mere classification logits. Our experimental results demonstrate that this alignment strategy is able to transfer the general audio knowledge to the specific domain of aero-engine bearing vibrations, enhancing the model’s ability to recognize and interpret complex fault patterns.
II-E Generative Fault Classification (GFC)
@@
GFC offers several substantial advantages over traditional classification approaches. It eliminates post-processing steps such as argmax operations and label mapping, which streamlines the diagnostic pipeline. The natural language outputs enhance interpretability by being immediately comprehensible to maintenance personnel without requiring specialized knowledge of model architectures or output interpretations. Additionally, the generative nature of GFC enables straightforward extensibility to new fault types without requiring architectural modifications, as the system can adapt to evolving aerospace industry simply by including examples of new faults in the training data. Experimental results demonstrate that GFC not only achieves superior classification accuracy compared to traditional approaches but also provides richer diagnostic information with practical utility for maintenance operations in aerospace applications.
III Experimental Validation
III-A Experimental Setup
@@
To systematically evaluate the performance of AeroGPT, a series of experiments were conducted using two widely-recognized bearing datasets: the aerospace bearing dataset from DIRG and the aero-engine bearing dataset from HIT. These datasets represent diverse operational conditions and fault characteristics relevant to aerospace applications, providing a robust foundation for validating our approach.
@@
The experimental validation encompasses three primary parts: First, an ablation study to assess the individual contributions of each component within the AeroGPT framework, specifically examining the impact of the Foundation Model (FM), Vibration Signal Alignment (VSA), and Generative Fault Classification (GFC). Second, comparative evaluations against state-of-the-art deep learning models to benchmark AeroGPT’s performance in terms of accuracy, precision, and F1-score. Thrid, qualitative analyses to demonstrate AeroGPT’s capability to generate interpretable, text-based diagnostic outputs that provide insights beyond mere fault classifications.
III-A1 Evaluation Metrics
@@
To quantitatively assess AeroGPT’s diagnostic performance, we employed standard classification metrics including accuracy, precision, and F1-score. These metrics provide complementary perspectives on the model’s effectiveness in correctly identifying bearing faults across diverse operational conditions:
$\text{Accuracy}=\frac{TP+TN}{TP+TN+FP+FN}$ (16)
$\text{Precision}=\frac{TP}{TP+FP}$ (17)
$\text{Recall}=\frac{TP}{TP+FN}$ (18)
@@
Here, TP (True Positive) represents correctly identified faulty conditions, TN (True Negative) denotes correctly identified healthy conditions, FP (False Positive) indicates healthy conditions incorrectly classified as faulty, and FN (False Negative) represents faulty conditions incorrectly classified as healthy. Additionally, we evaluated the model’s performance separately for non-defective and defective categories to assess its balanced capability across different fault conditions.
III-A2 Implementation Details
The implementation of AeroGPT was developed using the PyTorch framework, with the HuggingFace Transformers library utilized for model architecture and training utilities. All experiments were conducted on a single NVIDIA A10 GPU with 24GB of VRAM.
The foundation model used in the experiments was initialized on the weights of Qwen2-Audio [27], which comprises a pre-trained audio encoder with 124 million parameters and a large language model with 7 billion parameters. To fit the input of audio encoder, the vibration signals were sampled at 16kHz and segmented into wav files. For the Vibration Signal Alignment (VSA) stage, a dataset of 36,053 vibration-text paired samples was utilized to bridge the domain gap between general audio knowledge and bearing-specific vibration patterns. The datasets used for classification were split into training and test sets, with a split ratio of 8:2.
III-B Case 1: DIRG Bearing Dataset
III-B1 Dataset Description
The primary dataset for experimental validation was obtained from the Dynamic and Identification Research Group (DIRG) at Politecnico di Torino [28]. This aerospace-focused dataset features high-speed aeronautical bearings operating at speeds up to 35,000 rpm, accurately representing aerospace application conditions. Their bearing test rig comprises a shaft supported by three roller bearings, with controlled fault conditions introduced to one of them. Vibration data was collected using triaxial IEPE accelerometers mounted at two strategic locations on the bearing supports. The dataset encompasses seven distinct bearing conditions: one healthy state (0A) and six damage states (1A-6A) with precisely controlled conical indentations (150, 250, and 450 $\mu$m) on either the inner ring or a single roller. Data was recorded across multiple operational scenarios with shaft speeds ranging from 6,000 to 30,000 rpm and radial loads varying from 0 to 1800 N, at a sampling frequency of 51,200 Hz. To simulate the harsh operational environment, 4 dB of Gaussian white noise was added to the original signals. A total number of 22,134 training samples and 7,259 test samples were utilized.
@@ III-B1 Dataset Description
III-B2 Ablation Study
To comprehensively evaluate the contributions of each component in the proposed AeroGPT framework, we conducted an ablation study. TABLE I presents the performance metrics when including different combinations of the three primary components: Foundation Model (FM), which represents the model which has gone through Audio Knowledge Acquisition process, with general audio understanding capabilities; Vibration Signal Alignment (VSA), which bridges the domain gap between general audio and aero-engine bearing signal domain; and Generative Fault Classification (GFC), which employs fine-tuning techniques for direct fault label generation. All experiments maintained identical testing conditions with the same dataset split and evaluation metrics to ensure fair comparisons.
The results reveal that utilizing only the Foundation Model (FM) without any adaptation yields poor performance, with merely 14.87% accuracy, 6.31% precision, and 4.20% F1-score, which is approximately equivalent to random guessing. This underscores the significant domain gap between general audio understanding and aerospace bearing vibration analysis. When incorporating the VSA component while omitting GFC, a modest improvement can be observed in performance—accuracy increases by 5.78% to 20.65%, precision nearly doubles to 12.47%, and F1-score improves by 6.63% to 10.83%. This enhancement, while limited, indicates that the VSA stage establishes preliminary connections between general audio knowledge and vibration-specific patterns, though insufficient for accurate fault classification in isolation.
Notably, integrating GFC directly with the FM without the VSA intermediary yields a dramatic performance leap. The FM+GFC configuration achieves 97.21% accuracy, 97.84% precision and 97.52% F1-score, representing absolute improvements of 82.34%, 91.53%, and 93.32% in the respective metrics compared to using FM alone. This substantial enhancement demonstrates that PEFT enables effective adaptation of the foundation model to the target task, even without explicit domain alignment. The generative approach to fault classification proves highly effective in capturing the discriminative patterns in aerospace bearing vibration signals.
The complete AeroGPT framework, incorporating all three components (FM+VSA+GFC), achieves the highest performance across all metrics: 98.94% accuracy, 99.16% precision, and 99.02% F1-score. Compared to the FM+GFC configuration, the full framework yields further absolute improvements of 1.73%, 1.32%, and 1.50% in accuracy, precision, and F1-score, respectively. This corresponds to relative error reductions of 62.1% in misclassification rate, 61.1% in precision errors, and 60.5% in F1-score deficiency. These results confirm the synergistic effect of the VSA stage, which, while providing minimal benefits in isolation, significantly enhances the effectiveness of the subsequent GFC stage by aligning the foundation model’s representations with the aerospace bearing vibration signal domain.
The significant gap between configurations with and without GFC highlights the critical role of task-specific fine-tuning in adapting general-purpose models to specialized domains like aerospace bearing fault diagnosis. However, the performance improvement from adding VSA to GFC demonstrates that bridging the domain gap through intermediate pre-training remains valuable, especially for highly specialized technical applications such as aerospace propulsion systems monitoring.
III-B3 Comparison with Existing Methods
TABLE II: Performance comparison of AeroGPT against traditional deep learning models on the DIRG dataset.
@@ III-B3 Comparison with Existing Methods
AeroGPT 99.46% 97.08% 98.14% 98.42% 99.26% 99.07% 98.94% 99.16% 99.02%
@@
To comprehensively evaluate the efficacy of the proposed AeroGPT framework, we conducted comparative experiments against a diverse range of advanced deep learning architectures representing the current state-of-the-art in deep-learning-based classification tasks. The benchmark models span multiple architectural paradigms, including: MA1DCNN [13], a specialized multi-scale attention-enhanced convolutional architecture optimized for bearing vibration analysis; ResNet50 [29], a deep residual network that mitigates gradient vanishing through skip connections; ConvNeXt V2 [30], a modern convolutional design incorporating depthwise convolutions and layer normalization; Conv-Transformer, a hybrid architecture leveraging both local convolutional feature extraction and global attention mechanisms; Masked AutoEncoder [31], a self-supervised learning approach that reconstructs masked input segments; PoolFormer [32], which replaces attention mechanisms with simpler pooling operations while maintaining competitive performance; and the hierarchical Swin-Transformer architectures (original [33] and V2 [34]), which compute self-attention within shifted windows to achieve linear computational complexity with respect to image size. For vision-based architectures like ResNet and Swin-Transformer, vibration signals were converted to two-dimensional feature maps before processing.
As evidenced in TABLE II, AeroGPT achieves substantial performance advantages across nearly all evaluation metrics, with particularly pronounced improvements in defective state recognition. The overall accuracy of 98.94% represents a 2.73% improvement over the next best performer (MA1DCNN at 96.21%) and more than 8 percentage points above most transformer-based architectures. This performance gap underscores a fundamental limitation in conventional approaches: despite their sophisticated architectures, models designed for generic feature extraction struggle to capture the patterns of bearing faults without domain-specific adaptation.
A distinctive pattern emerges when examining performance across defective and non-defective categories. Traditional architectures exhibit a substantial performance asymmetry, with accuracy discrepancies exceeding 20 percentage points in several cases (e.g., ConvNeXt V2: 95.55% vs. 72.05%; Conv-Transformer: 96.11% vs. 71.16%; PoolFormer: 94.79% vs. 67.19%). This indicates a fundamental challenge in discriminating between different fault types compared to distinguishing between healthy and faulty states. In contrast, AeroGPT demonstrates remarkable category balance, with comparable performance across both non-defective (99.46%) and defective (98.42%) classifications. This balanced capability is particularly critical in aerospace applications, where missed defects can lead to catastrophic failures and false alarms result in unnecessary and costly maintenance interventions.
The performance advantage of AeroGPT stems from its architectural advantages. First, unlike conventional models that build representations from scratch, AeroGPT leverages and adapts the rich acoustic domain knowledge, which provides a sophisticated basis for interpreting mechanical vibration patterns that share fundamental properties with acoustic phenomena. Second, the VSA stage effectively bridges the domain gap between general audio and aerospace-specific vibration patterns, as confirmed by the ablation study results.
Notably, even the most advanced transformer-based architectures like Swin-Transformer V2, which achieve state-of-the-art performance in many computer vision tasks, fall significantly short in this domain-specific application. This emphasizes that architectural sophistication alone is insufficient for specialized engineering tasks—domain knowledge integration.
III-B4 Qualitative Analysis

### v2

@@ III-B Audio Knowledge Acquisition
The final refinement stage implements Direct Preference Optimization (DPO), where alternative responses to the same audio input are evaluated through a comparative scoring mechanism. The optimization objective is formulated as:
@@
where $\mathcal{P}_{\theta}$ represents the system being optimized, $\mathcal{P}_{ref}$ is a reference system, $y_{w}$ and $y_{l}$ are the preferred and less preferred responses respectively, $\sigma$ is the sigmoid function, and $\beta$ is a hyperparameter controlling optimization strength. This multi-stage process produces a large-scale audio foundation model with sophisticated acoustic pattern recognition capabilities, providing the foundation for interpreting bearing vibration signals as acoustic phenomena.
III-C LoRA-based Domain Knowledge Adaptation
@@
The adaptation process operates on the principle that domain-specific knowledge can be effectively integrated through low-rank decomposition of weight updates. Rather than modifying the entire weight matrices of the large-scale model, which would require enormous computational resources, AeroGPT freezes the pre-trained weight matrices $\mathbf{W}_{0}\in\mathbb{R}^{d\times k}$ and introduces trainable low-rank adaptations through:
@@
where $\mathbf{B}\in\mathbb{R}^{d\times r}$ and $\mathbf{A}\in\mathbb{R}^{r\times k}$ are low-rank matrices with rank $r\ll\min(d,k)$, $\mathbf{x}$ is the input, $\mathbf{h}$ is the output, and $\alpha$ is a scaling factor for training stability. This factorization significantly reduces trainable parameters from $d\times k$ to $r\times(d+k)$, achieving over 99% parameter reduction and reducing VRAM usage from over 100 GB for full fine-tuning to approximately 17 GB.
In our framework, LoRA adaptations are applied to all linear layers within the audio encoder, aligner, and LLM components, including self-attention mechanisms (query, key, value projections) and feed-forward networks. The LoRA rank $r$ is set to 16 and the scaling factor $\alpha$ is set to 32. Matrix $\mathbf{A}$ is initialized with random Gaussian values, while $\mathbf{B}$ is initialized with zeros to ensure $\Delta\mathbf{W}=\mathbf{B}\mathbf{A}=0$ at training start. Through this adaptation process, AeroGPT bridges the gap between general audio understanding and aero-engine bearing vibrations.
III-D Vibration Signal Alignment (VSA)
The Vibration Signal Alignment (VSA) mechanism is designed to bridge the fundamental gap between general audio knowledge and the domain-specific characteristics of aero-engine bearing vibrations. This is achieved through an alignment process that preserves the rich hierarchical representations inherent in large-scale audio models while adapting them to the specialized patterns of mechanical vibrations.
@@ III-D Vibration Signal Alignment (VSA)
This training paradigm enables the model to learn contextual relationships between vibration patterns and diagnostic interpretations in an end-to-end manner while preserving the linguistic capabilities of the underlying large language model, allowing for the generation of interpretable diagnostic outputs rather than mere classification logits. Our experimental results demonstrate that this alignment strategy is able to transfer the general audio knowledge to the specific domain of aero-engine bearing vibrations, enhancing the model’s ability to recognize and interpret complex fault patterns.
The VSA stage utilizes the Paderborn University bearing dataset [17] for constructing vibration-text pairs. This dataset contains 32 bearing samples across three fault categories (healthy, inner ring damage, outer ring damage) with diverse damage mechanisms: 6 healthy bearings with different run-in periods, 12 artificially damaged bearings created through EDM machining, drilling, and electric engraving, and 14 real-damage bearings exhibiting fatigue pitting and plastic deformation from accelerated life tests. We use this dataset for VSA rather than benchmarking because while its coarse-grained labels with only inner/outer ring distinction are too simple for fine-grained evaluation, its diverse damage mechanisms are ideal for learning generalizable vibration-semantic associations.
Since publicly available bearing fault datasets provide only numerical class labels without natural language descriptions, we constructed a dataset of 36,053 vibration-text pairs using a semi-automatic method that combines expression templates with domain knowledge. Per-sample human authoring was avoided due to extreme effort and the risk of inconsistent phrasing in large-scale corpora. The design objective was to keep each description faithful to the physical state of the signal while enabling wording diversity.
[…]
Natural language descriptions were then generated by filling several sentence templates with these fields. The templates employ different sentence structures and phrase variants (e.g., “inner race”/“inner ring”, “impact pulses”/“impulsive bursts”). During generation, we randomly select a template and phrase variants, yielding descriptions that remain faithful to the underlying data while adding modest variety. This process provides consistent supervision for VSA and improves robustness and generalization.
It is important to note that template-based generation has inherent limitations. The linguistic diversity is constrained by the number of templates and synonym substitutions, which may cause the model to learn “template style” rather than fully open-ended diagnostic language. However, we try to mitigate these concerns through employing multiple sentence structures and safe phrase variants to introduce linguistic variety, including descriptive content about signal characteristics linked to fault types based on established diagnostic patterns, and monitoring throughout the training process to avoid overfitting to the templates and forgetting the true semantic meaning.
III-E Generative Fault Classification (GFC)
@@
GFC offers several substantial advantages over traditional classification approaches. It eliminates post-processing steps such as argmax operations and label mapping, which streamlines the diagnostic pipeline. The natural language outputs enhance interpretability by being immediately comprehensible to maintenance personnel without requiring specialized knowledge of model architectures or output interpretations. Additionally, the generative nature of GFC enables straightforward extensibility to new fault types without requiring architectural modifications, as the system can adapt to the evolving aerospace industry simply by including examples of new faults in the training data. Experimental results demonstrate that GFC not only achieves superior classification accuracy compared to traditional approaches but also provides richer diagnostic information with practical utility for maintenance operations in aerospace applications.
IV Experimental Validation
IV-A Experimental Setup
@@
To systematically evaluate the performance of AeroGPT, a series of experiments were conducted using two widely recognized bearing datasets: the aerospace bearing dataset from DIRG and the aero-engine bearing dataset from HIT. These datasets represent diverse operational conditions and fault characteristics relevant to aerospace applications, providing a robust foundation for validating our approach.
@@
The experimental validation encompasses three primary parts: First, an ablation study to assess the individual contributions of each component within the AeroGPT framework, specifically examining the impact of the Foundation Model (FM), Vibration Signal Alignment (VSA), and Generative Fault Classification (GFC). Second, comparative evaluations against state-of-the-art deep learning models to benchmark AeroGPT’s performance in terms of accuracy, precision, and F1-score. Third, qualitative analyses to demonstrate AeroGPT’s capability to generate interpretable, text-based diagnostic outputs that provide insights beyond mere fault classifications.
IV-A1 Evaluation Metrics
@@
To quantitatively assess AeroGPT’s diagnostic performance, we employed standard classification metrics including accuracy, precision, and F1-score. These metrics provide complementary perspectives on the model’s effectiveness in correctly identifying bearing faults across diverse operational conditions. Additionally, we evaluated the model’s performance separately for non-defective and defective categories to assess its balanced capability across different fault conditions.
IV-A2 Implementation Details
AeroGPT was implemented using the PyTorch framework, with the HuggingFace Transformers library utilized for model architecture and training utilities. All experiments were conducted on a single NVIDIA A10 GPU with 24 GB of VRAM.
The foundation model used in the experiments was initialized with the weights of Qwen2-Audio [5], which comprises a pre-trained audio encoder with 124 million parameters and a large language model with 7 billion parameters. To prepare vibration signals for input to the audio encoder, several signal processing steps were applied. First, all vibration signals underwent resampling to standardize the sampling rate to 16 kHz, ensuring compatibility with the pre-trained audio encoder. Second, amplitude normalization was performed on each signal segment to eliminate variations in signal strength across different operating conditions and sensor configurations. Specifically, peak normalization was applied to linearly scale each segment’s amplitude to a fixed range of [-1.0, 1.0], enabling the model to focus on structural patterns and waveform characteristics rather than absolute signal intensity. Third, the normalized signals were quantized to 16-bit Pulse-Code Modulation (PCM) format and saved as WAV (.wav) files, maintaining the standard audio format expected by the encoder while preserving signal fidelity. For the Vibration Signal Alignment (VSA) stage, a dataset of 36,053 vibration-text paired samples was utilized to bridge the domain gap between general audio knowledge and bearing-specific vibration patterns. The datasets used for classification were split into training and test sets, with a split ratio of 8:2.
IV-B Case 1: DIRG Bearing Dataset
IV-B1 Dataset Description
@@
IV-B2 Ablation Study
TABLE I presents the ablation study results for three components: Foundation Model (FM) with general audio understanding, Vibration Signal Alignment (VSA) for domain adaptation, and Generative Fault Classification (GFC) for direct label generation. For FM+VSA, evaluation adopts a zero-shot protocol where the model selects from predefined fault categories, with top-1 accuracy computed based on exact matches.
Using FM alone yields only 14.87% accuracy, approximately equivalent to random guessing, underscoring the significant domain gap. Adding VSA improves accuracy to 20.65%, indicating preliminary connections between audio knowledge and vibration patterns. However, FM+GFC achieves 97.21% accuracy, demonstrating that task-specific fine-tuning enables effective adaptation even without explicit domain alignment.
The complete framework (FM+VSA+GFC) achieves the highest performance: 98.94% accuracy, 99.16% precision, and 99.02% F1-score, corresponding to relative error reductions of 62.1%, 61.1%, and 60.5% compared to FM+GFC. These results confirm the synergistic effect of VSA, which, while providing minimal benefits in isolation, significantly enhances GFC effectiveness by aligning representations with the aerospace bearing vibration domain.
IV-B3 Comparison with Existing Methods
TABLE II: Performance comparison of AeroGPT against both fault diagnosis models and general-purpose deep learning models on the DIRG dataset.
@@ IV-B3 Comparison with Existing Methods
Fault Diagnosis Models
AeroGPT (Ours) 99.46% 97.08% 98.14% 98.42% 99.26% 99.07% 98.94% 99.16% 99.02%
DCNDSC 99.20% 98.83% 97.21% 96.14% 97.48% 97.75% 97.67% 97.71% 97.68%
LiConvFormer 98.87% 97.08% 95.97% 93.57% 96.08% 96.26% 96.22% 96.24% 96.22%
TFN-STFF 96.90% 87.06% 89.46% 88.20% 93.53% 93.08% 92.55% 92.60% 92.55%
@@
ILDM 96.35% 84.10% 87.78% 76.07% 86.60% 85.93% 86.21% 86.37% 86.19%
MPSO-ACBCNN 99.01% 95.73% 96.56% 95.67% 97.61% 97.47% 97.34% 97.35% 97.34%
TAUN 99.04% 98.21% 96.57% 93.50% 95.96% 96.22% 96.27% 96.30% 96.28%
DSRSN 97.51% 86.77% 91.78% 91.56% 96.01% 95.02% 94.53% 94.77% 94.56%
SPCFormer 98.10% 93.68% 93.32% 84.90% 91.14% 91.20% 91.50% 91.61% 91.54%
MRCNN-LSTM 93.98% 78.63% 79.04% 68.30% 81.56% 81.49% 81.14% 81.25% 81.17%
General Purpose Models
@@
To comprehensively assess AeroGPT’s effectiveness, we benchmarked it against two categories of models: specialized fault diagnosis methods and general-purpose deep learning architectures. The fault diagnosis baselines include DCNDSC [32], a large-scale dense connectivity framework employing depthwise separable convolutions for multi-machine diagnostics; LiConvFormer [47], which combines separable multiscale convolutions with broadcast self-attention; TFN (with STFF) [3], a time-frequency network for bearing fault detection; MA1DCNN [40], a multi-scale attention convolutional network for vibration analysis; ILDM [14], a large-scale diagnostic model approach utilizing time-series embeddings with 1D-2D-1D transformations; DSRSN [46], a deep spiking residual shrinkage network that introduces attention mechanisms and soft thresholding to improve recognition under high-noise conditions; TAUN [16], a traceable algorithm unrolling network that constructs an interpretable feature extractor by unrolling the iterative sparse coding algorithm; MPSO-ACBCNN [51], an automated CNN design method using modified particle swarm optimization with advanced convolution blocks for satellite attitude control system fault diagnosis; MRCNN-LSTM [25], a model fusion approach combining multiscale residual CNN with LSTM to extract both spatial and temporal features; and SPCFormer [9], a lightweight Transformer with selective patches and channels modules. The general-purpose models comprise ResNet50 [11], ConvNeXt V2 [44], Conv-Transformer, Masked AutoEncoder [10], PoolFormer [48], and Swin-Transformer variants [27, 26]. For vision-based architectures, 1D vibration signals were reshaped into approximately square 2D feature maps of size $d\times d$ where $d=\lfloor\sqrt{L}\rfloor$, with zero-padding applied when necessary.
As shown in TABLE II, AeroGPT achieves superior performance across virtually all metrics, with overall accuracy reaching 98.94%. Among fault diagnosis baselines, DCNDSC (97.67%) and MPSO-ACBCNN (97.34%) demonstrate strong performance through their specialized architectural designs, while TAUN (96.27%) and LiConvFormer (96.22%) also achieve competitive accuracy. In contrast, methods such as SPCFormer (91.50%) and MRCNN-LSTM (81.14%) show limited effectiveness on this dataset, likely due to their original design focus on different application domains. General-purpose models exhibit even larger performance gaps, with ResNet50 achieving only 92.75% accuracy. These results demonstrate that domain-agnostic architectures, despite their sophistication, cannot adequately capture fault-specific patterns without specialized adaptation.
A critical observation concerns the performance disparity between defective and non-defective states. General-purpose models exhibit severe imbalance, with accuracy gaps exceeding 20 percentage points (e.g., ConvNeXt V2: 95.55% vs. 72.05%; PoolFormer: 94.79% vs. 67.19%). Even specialized models like TFN-STFF (96.90% vs. 88.20%) and ILDM (96.35% vs. 76.07%) show notable degradation on defective samples, indicating difficulties in discriminating among diverse fault types. Conversely, AeroGPT maintains exceptional balance with only a 1.04 percentage difference (99.46% vs. 98.42%), which is crucial for aerospace applications where both missed detections and false alarms carry substantial operational consequences.
AeroGPT’s advantages stem from two key factors. First, it leverages pre-trained acoustic representations that inherently capture temporal-spectral patterns relevant to mechanical vibrations, rather than learning from scratch. Second, the VSA adaptation stage effectively transfers this knowledge to aerospace-specific fault characteristics, as validated by ablation studies.
IV-B4 Qualitative Analysis